# IFJPAN-IV-2023-3

### v1

@@ ABSTRACT
Anomalous contributions to the electric and magnetic dipole moments of the $\tau$ lepton from new physics scenarios have brought renewed interest in the development of new charge-parity violating signatures in $\tau$ pair production at Belle II energies, and also at higher energies of the Large Hadron Collider and the Future Circular Collider. In this paper, we discuss the effects of spin correlations, including transverse degrees of freedom, in the $\tau$ pair production and decay. These studies include calculating analytical formulas, obtaining numerical results, and building semi-realistic observables sensitive to the transverse spin correlations induced by the dipole moments of the $\tau$ lepton.
The effects of such anomalous contributions to the dipole moments are introduced on top of precision simulations of $e^{-}e^{+}\to\tau^{-}\tau^{+}$, $q\bar{q}\to\tau^{-}\tau^{+}$ and $\gamma\gamma\to\tau^{-}\tau^{+}$ processes, involving multi-body final states. Respective extensions of the Standard Model amplitudes and the reweighting algorithms are implemented into the KKMC Monte Carlo, which is used to simulate $\tau$ pair production in $e^{-}e^{+}$ collisions, and the TauSpinner program, which is used to reweight events with $\tau$ pair produced in $pp$ collisions.
@@ 1 Introduction
The KKMC Monte Carlo has been used to generate events of the SM $e^{-}e^{+}\to\tau^{-}\tau^{+}$ process for several decades, and is now enriched with several additional processes to describe NP effects potentially observable at $e^{-}e^{+}$ collider and the Belle II experiment [10]. In particular, recent development in Ref. [11] allows one to include anomalous contributions to electric and magnetic dipole moments of the $\tau$ lepton, in case of low energy $\tau$-pair production. In KKMC, all spin correlations are rigorously implemented, including quantum entanglement effects. This is the case for configurations where an arbitrary number of hard bremsstrahlung photons can be present. It can also provide valuable benchmarks for spin effects in $e^{-}e^{+}\to\tau^{-}\tau^{+}$ process, with $\tau$ decays included.
The TauSpinner program is a convenient tool to study observables sensitive to the NP effects at hadron colliders. It allows one to include NP and spin effects, which are absent in event samples generated with general purpose MC generators like Pythia [12] and Sherpa [13]. At the beginning, the longitudinal spin effects in the case of the Drell-Yan processes and the production of $\tau$ leptons in Higgs decays at the Large Hadron Collider (LHC) were implemented [14]. Later, they were extended to applications of the NP interactions [15] in the hard processes, in which lepton pairs are accompanied with one or two hard jets [15], and to include the transverse spin effects [16]. The list of applications was also extended to incorporate the electroweak effects [17, 18].
@@
In this paper, we focus on the effects of dipole moments/form-factors on spin correlations in the $\tau$ pair production and decay. Quantifying those effects requires explicit calculation of the matrix elements including longitudinal and transverse spin components and their correlations. First, we consider the SM case, because it determines the nature of interfering contributions for analysis of anomalous dipole moments due to NP. In the second step, we consider the anomalous dipole moments themselves. These studies include calculation of the analytical formulas, validations of the tools, and an attempt to build semi-realistic observables sensitive to transverse spin correlations induced by the dipole moments. Impact on the size of cross sections for $\tau$ pair production is addressed marginally, as it was already discussed in the literature [25].
Technically, the effect of anomalous dipole moments can be introduced on top of precision simulations of $e^{-}e^{+}\to\tau^{-}\tau^{+}$, $q\bar{q}\to\tau^{-}\tau^{+}$ and $\gamma\gamma\to\tau^{-}\tau^{+}$ processes, involving multi-body final states and correlations between $\tau$ decay products. The Monte Carlo techniques using event-by-event weights, are convenient for this purpose. The KKMC Monte Carlo is used for simulating spin-correlation effects in $\tau$ pair produced in $e^{-}e^{+}$ collisions, while the TauSpinner program is used for reweighting events where $\tau$ pairs are produced in $pp$ collisions.
Our paper is organized in the following manner. In Section 2, we present analytic calculation of the spin correlation matrix, which include effects of the $\tau$-lepton anomalous magnetic and electric dipole form-factors. Analytic formulas, described in Section 3, are useful for implementation of transverse spin correlations for each of the parton level processes $q\bar{q}\to\tau^{-}\tau^{+}$, $e^{-}e^{+}\to\tau^{-}\tau^{+}$ and $\gamma\gamma\to\tau^{-}\tau^{+}$. In Sections 4, we recall main points of the KKMC and TauSpinner reweighting algorithms for inclusion of the dipole moments. Discussions on the possible test observables and some numerical results are collected in Section 5. Section 6 closes the paper with a summary and outlook.
@@ 2 Amplitudes and spin correlations
In Ref. [11], formulas for including anomalous magnetic and electric dipole form-factors in elementary $2\to 2$ parton processes of $\tau$-pair production were discussed. The elementary parton-level and improved Born-level formulas were presented there for $e^{-}e^{+}\to\tau^{-}\tau^{+}$ process, including details of reference frames [^1] A comment on the direction of reference frames is in place here. Formally speaking, the interchange of reaction name between $e^{-}e^{+}\to\tau^{-}\tau^{+}$ and $e^{+}e^{-}\to\tau^{+}\tau^{-}$ is purely conventional and not of physics content. The choice however does imply which among the incoming $e^{-}$ or $e^{+}$ particle is taken as the first beam in the positive $\hat{z}$ direction. The second choice was used in some of older projects. Nowdays, the choice of $e^{-}$ is made more frequently, as the more energetic beam in asymmetric colliders. Due to overlapping conventions, a sign mismatch may occur. In our programs, a rotation by an angle of $\pi$ around the $\hat{y}$ axis may be necessary before the anomalous contributions are inserted into the other parts of the code. used in spin-correlation matrix calculation. Those formulas are however limited to low energies, where contribution of $Z$-boson exchange can be neglected in specific applications.
For higher energy $e^{-}e^{+}\to\tau^{-}\tau^{+}$, and also for $q\bar{q}\to\tau^{-}\tau^{+}$ parton level processes at the LHC, inclusion of the contribution due to the exchange of $Z$-boson is necessary. For very high $\tau$-pair virtuality of more than 160160~{}160160 GeV, the contributions from doubly resonant $WW$ and $ZZ$ boxes need to be taken into account as well, which however will not be covered in this paper. Here we extend results of Ref. [11], and derive the formulas including $Z$ and $\gamma^{*}$ exchanges and their interference, as applicable for a virtuality of the $\tau$-pair upto and slightly above the $Z$-peak.
To complete the picture we also derive the formulas for $\gamma\gamma\to\tau^{-}\tau^{+}$ processes in the so-called light-by-light scattering approximation, which recently became of high interest in the analysis of heavy ion beams collisions at the LHC [5, 6].
@@
In the center-of-mass (CM) frame, the components of the momenta are
@@ 2.1 The $f_{i}\bar{f}_{i}\to\tau^{-}\tau^{+}$, $i=e,u,d$ case
$\Gamma^{\mu}_{\gamma}(q)=e\,\Big{\{}\gamma^{\mu}+\frac{\sigma^{\mu\nu}q_{\nu}}{2m}\,\big{[}iA(s)+B(s)\gamma_{5}\big{]}\Big{\}},$ (3)
where $\sigma^{\mu\nu}=\frac{i}{2}[\gamma^{\mu},\,\gamma^{\nu}]$, $e$ is the charge of the positron, $A(s)=F_{2}(q^{2})$ is the Pauli form-factor, and $B(s)=F_{3}(q^{2})$ is the electric dipole form-factor, which depend on $s=q^{2}$, where $q=k_{1}+k_{2}=p_{-}+p_{+}$. At the on-shell photon point, $A(0)$ is an anomalous dipole moment $a$, while $B(0)$ is related to the $CP$ violating electric dipole moment $d$,
$A(0)=a=\frac{1}{2}(g-2),\qquad\quad B(0)=\frac{2m_{\tau}}{e}d,$ (4)
where $g$ is the gyro-magnetic factor.
@@
$\Gamma^{\mu}_{Z}(q)=\frac{g_{Z}}{2}\,\Big{\{}\gamma^{\mu}(v_{\tau}-\gamma_{5}a_{\tau})+\frac{\sigma^{\mu\nu}q_{\nu}}{2m}\,\big{[}iX(s)+Y(s)\gamma_{5}\big{]}\Big{\}},$ (7)
[…]
@@
$u$ $+\frac{2}{3}$ $+\frac{1}{2}-\frac{4}{3}\sin^{2}\theta_{W}$ $+\frac{1}{2}$
@@
$d$ $-\frac{1}{3}$ $-\frac{1}{2}+\frac{2}{3}\sin^{2}\theta_{W}$ $-\frac{1}{2}$
@@
To simplify calculations, it is convenient to use the Gordon identities for the matrix elements of $\Gamma_{\gamma}^{\mu}(q)$ and $\Gamma_{Z}^{\mu}(q)$. This allows one to rewrite the respective currents as:
$\bar{u}(p_{-})\,\Gamma^{\mu}_{\gamma}(q)\,v(p_{+})$ $=$ $e\,\bar{u}(p_{-})\Big{\{}\gamma^{\mu}(1+A(s))$ (8)
$+\frac{(p_{+}-p_{-})^{\mu}}{2m}\big{[}A(s)-iB(s)\gamma_{5}\big{]}\Big{\}}v(p_{+}),$
$\bar{u}(p_{-})\,\Gamma^{\mu}_{Z}(q)\,v(p_{+})$ $=$ $\frac{g_{Z}}{2}\,\bar{u}(p_{-})\Big{\{}\gamma^{\mu}((v_{\tau}+X(s))-\gamma_{5}a_{\tau})$ (9)
$+\frac{(p_{+}-p_{-})^{\mu}}{2m}\big{[}X(s)-iY(s)\gamma_{5}\big{]}\Big{\}}v(p_{+}).$
@@
If $\tau$ decays are taken into account, the vectors defining $\tau^{-}$ and $\tau^{+}$ density matrices, $s^{-}_{i}$ and $s^{+}_{j}$ in Eq. (14), are replaced respectively by the polarimetric vectors, depending on the $\tau$ decay matrix elements, $h^{-}_{i}$ and $h^{+}_{j}$.
@@
where $R^{(\gamma)}_{i,j}$, $R^{(Z)}_{i,j}$ and $R^{(\gamma Z)}_{i,j}$ represent contributions due to $\gamma$ exchange, $Z$-boson exchange and $\gamma\,Z$ interference, respectively. In the following, we include only terms linear in all form-factors. The quantization of $\tau^{\mp}$ are performed in their respective rest frames. These frames differ from the reaction frame, defined by Eq. (2), by the boosts along $\tau^{\mp}$ directions.
@@
$R_{11}^{(Z)}=\frac{g_{Z}^{4}m^{4}\gamma^{2}}{4N_{i}D_{Z}(s)}\sin^{2}(\theta)\big{\{}\left(1-\gamma^{2}\right)a_{\tau}^{2}+v_{\tau}\bigl{[}4\gamma^{2}\text{Re}(X(s))$
@@
$R_{12}^{(Z)}=-\frac{g_{Z}^{4}m^{4}\beta\gamma^{4}}{2N_{i}D_{Z}(s)}{\sin^{2}(\theta)\big{[}\text{Im}(X(s))a_{\tau}-\text{Re}(Y(s))v_{\tau}\big{]}\left(a_{i}^{2}+v_{i}^{2}\right)},$
$R_{22}^{(Z)}=\frac{g_{Z}^{4}m^{4}\gamma^{2}}{4N_{i}D_{Z}(s)}\left(\gamma^{2}-1\right){\sin}^{2}(\theta)\left(a_{\tau}^{2}-v_{\tau}^{2}\right)\left(a_{i}^{2}+v_{i}^{2}\right),$
$R_{13}^{(Z)}=\frac{g_{Z}^{4}m^{4}\gamma^{3}}{4N_{i}D_{Z}(s)}\big{\{}4\,\beta\,\text{sin}(\theta)\,a_{i}\left[\gamma^{2}\text{Re}(X(s))a_{\tau}+\left(-\gamma^{2}\text{Im}(Y(s))+a_{\tau}\right)v_{\tau}\right]v_{i}$
@@
$R_{23}^{(Z)}=-\frac{g_{Z}^{4}m^{4}\gamma^{3}}{2N_{i}D_{Z}(s)}\text{sin}(\theta)\big{\{}\beta\gamma^{2}\text{Re}(Y(s))\left[v_{\tau}\text{cos}(\theta)(a_{i}^{2}+v_{i}^{2})+2\beta a_{\tau}a_{i}v_{i}\right]$

### v2

@@ ABSTRACT
Anomalous contributions to the electric and magnetic dipole moments of the $\tau$ lepton from new physics scenarios have brought renewed interest in the development of new charge-parity violating signatures in $\tau$-pair production at Belle II energies, and also at higher energies of the Large Hadron Collider and the Future Circular Collider. In this paper, we discuss the effects of spin correlations, including transverse degrees of freedom, in the $\tau$-pair production and decay. These studies include calculating analytical formulas, obtaining numerical results, and building semi-realistic observables sensitive to the transverse spin correlations induced by the dipole moments of the $\tau$ lepton. The effects of such anomalous contributions to the dipole moments are introduced on top of precision simulations of $e^{-}e^{+}\to\tau^{-}\tau^{+}$, $q\bar{q}\to\tau^{-}\tau^{+}$ and $\gamma\gamma\to\tau^{-}\tau^{+}$ processes, involving multi-body final states. The $\tau$ decays are simulated along with radiative corrections, in particular electroweak box contributions of $WW$ and $ZZ$ exchanges are taken into account. Respective extensions of the Standard Model amplitudes and the reweighting algorithms are implemented into the KKMC Monte Carlo, which is used to simulate $\tau$-pair production in $e^{-}e^{+}$ collisions, and the TauSpinner program, which is used to reweight events with $\tau$ pair produced in $pp$ collisions.
@@ 1 Introduction
The KKMC Monte Carlo (MC) has been used to generate events of the SM $e^{-}e^{+}\to\tau^{-}\tau^{+}$ process for several decades, and is now enriched with several additional processes to describe NP effects potentially observable at $e^{-}e^{+}$ collider and the Belle II experiment [10]. In particular, recent development in Ref. [11] allows one to include anomalous contributions to electric and magnetic dipole moments of the $\tau$ lepton, in case of low energy $\tau$-pair production. In KKMC, all spin correlations are rigorously implemented, including quantum entanglement effects. This is the case for configurations where an arbitrary number of hard bremsstrahlung photons can be present. It can also provide valuable benchmarks for spin effects in $e^{-}e^{+}\to\tau^{-}\tau^{+}$ process, with $\tau$ decays included.
The TauSpinner program is a convenient tool to study and prepare observables sensitive to the NP effects at hadron colliders. Its purpose is to introduce, with the help of weights, small effects on top of event samples of the SM content. It is imperative, as in case of KKMC, to evaluate if the precision of the SM simulation is sufficient for the user’s needs. Then, TauSpinner may be used to introduce NP and/or spin effects. The samples may be generated with the help of general purpose MC generators like Pythia [12], Sherpa [13] or with the help of other, sophisticated generator setup. In its earlier versions, TauSpinner was used to introduce the longitudinal spin effects, in the case of the Drell-Yan processes or the Higgs production with decays into $\tau$ leptons, at the Large Hadron Collider (LHC) [14]. Later, they were extended to applications of the NP interactions [15] in the hard processes, in which lepton pairs are accompanied with one or two hard jets [15], and to include the transverse spin effects [16]. The list of applications was also extended to incorporate the electroweak (EW) effects [17, 18].
@@
In this paper, we focus on the effects of dipole moments/form-factors on spin correlations in the $\tau$-pair production and decay. Quantifying those effects requires explicit calculation of the matrix elements including longitudinal and transverse spin components and their correlations.
First, we consider the SM case, because it determines the nature of interfering contributions for analysis of anomalous dipole moments due to NP. The SM amplitude is evaluated in the improved Born approximation (IBA). The theoretical basis of IBA is formulated in Ref. [25], and our approach is explained in Ref. [17]. It allows us to separate the EW correction into pure QED part and the remaining one, sometimes called the genuine weak part. However, we shorten the name to simply refer to this as the EW correction. Numerically these corrections are included in IBA and rely on the EW Dizet library [25], which is installed in KKMC. The EW corrections are introduced with form-factor corrections to the SM couplings and propagators, which enter the IBA amplitude used for calculation of EW weights. They represent complete ${\cal O}(\alpha)$ EW corrections with QED contributions removed but augmented with carefully selected dominant higher-order terms. The $WW$ and $ZZ$ box-diagram contributions, which are numerically important above the $WW$ and $ZZ$ threshold, are thus accounted for. This approach was very successful in analyses of LEP I precision physics, and later extended for use in Tevatron and LHC precision physics.
In the second step, we consider the anomalous dipole moments themselves. These studies include calculation of the analytical formulas, validations of the tools, and an attempt to build semi-realistic observables sensitive to transverse spin correlations induced by the dipole moments. Impact on the size of cross sections for $\tau$-pair production is addressed marginally. Note, that effect of dipole moments in the radiative decay $Z\to\tau^{-}\tau^{+}\gamma$ based on LEP data was discussed in Ref. [26], and for $\gamma\gamma\to\tau^{-}\tau^{+}$ in peripheral Pb+Pb collisions in Ref. [27].
Technically, the effect of anomalous dipole moments can be introduced on top of precision simulations of $e^{-}e^{+}\to\tau^{-}\tau^{+}$, $q\bar{q}\to\tau^{-}\tau^{+}$ and $\gamma\gamma\to\tau^{-}\tau^{+}$ processes, involving multi-body final states and correlations between $\tau$ decay products. The MC techniques using event-by-event weights, are convenient for this purpose. The KKMC MC is used for simulating spin-correlation effects in $\tau$ pair produced in $e^{-}e^{+}$ collisions, while the TauSpinner program is used for reweighting events where $\tau$ pairs are produced in $pp$ collisions.
Our paper is organized in the following manner. In Sec. 2, we present analytic calculation of the spin correlation matrix, which include effects of the $\tau$-lepton anomalous magnetic and electric dipole form-factors. Analytic formulas, described in Sec. 3, are useful for implementation of transverse spin correlations for each of the parton level processes $q\bar{q}\to\tau^{-}\tau^{+}$, $e^{-}e^{+}\to\tau^{-}\tau^{+}$ and $\gamma\gamma\to\tau^{-}\tau^{+}$. In Sec. 4, we recall main points of the KKMC and TauSpinner reweighting algorithms for inclusion of the dipole moments. Discussions on the possible test observables and some numerical results are collected in Sec. 5. Sec. 6 closes the paper with a summary and outlook.
@@ 2 Amplitudes and spin correlations
In Ref. [11], formulas for including anomalous magnetic and electric dipole form-factors in elementary $2\to 2$ parton processes of $\tau$-pair production were discussed. The elementary parton-level and improved Born-level formulas were presented there for $e^{-}e^{+}\to\tau^{-}\tau^{+}$ process, including details of reference frames [^1] A comment on the direction of reference frames is in place here. Formally speaking, the interchange of reaction name between $e^{-}e^{+}\to\tau^{-}\tau^{+}$ and $e^{+}e^{-}\to\tau^{+}\tau^{-}$ is purely conventional and not of physics content. The choice however does imply which among the incoming $e^{-}$ or $e^{+}$ particle is taken as the first beam (analogously if $\tau^{-}$ or $\tau^{+}$ is taken as the first outgoing lepton) in the positive $\hat{z}$ direction. The second choice was used in some of older projects. Nowadays, the choice of $e^{-}$ is made more frequently, as the more energetic beam in asymmetric colliders. Due to overlapping conventions, a sign mismatch may occur. In our programs, a rotation by an angle of $\pi$ around the $\hat{y}$ axis may be necessary before the anomalous contributions are inserted into the other parts of the code. used in spin-correlation matrix calculation. Those formulas are however limited to low energies, where contribution of $Z$-boson exchange can be neglected in specific applications.
For higher energy $e^{-}e^{+}\to\tau^{-}\tau^{+}$, and also for $q\bar{q}\to\tau^{-}\tau^{+}$ parton level processes at the LHC, inclusion of the contribution due to the exchange of $Z$-boson is necessary. For very high $\tau$-pair virtuality of more than 160160~{}160160 GeV, the contributions from doubly resonant $WW$ and $ZZ$ boxes need to be taken into account as well, following for example, Refs. [17, 28] on TauSpinner and KKMC, where EW loop corrections installation is documented. Here we extend results of Ref. [11], and derive the formulas including $Z$ and $\gamma$ exchanges and their interference.
To complete the picture we also derive the formulas for $\gamma\gamma\to\tau^{-}\tau^{+}$ processes in the so-called light-by-light scattering approximation, which recently became of high interest in the analysis of heavy-ion beams collisions at the LHC [5, 6].
@@
In the center-of-mass (c.m.) frame, the components of the momenta are
@@ 2.1 The $f_{i}\bar{f}_{i}\to\tau^{-}\tau^{+}$, $i=e,u,d$ case
Dipole form-factors and improved Born approximation
@@
$\Gamma^{\mu}_{\gamma}(q)=-ie{{Q_{\tau}}}\Big{\{}\gamma^{\mu}+\frac{\sigma^{\mu\nu}q_{\nu}}{2m_{\tau}}\,\big{[}iA(s)+B(s)\gamma_{5}\big{]}\Big{\}},$ (3)
where $\sigma^{\mu\nu}=\frac{i}{2}[\gamma^{\mu},\,\gamma^{\nu}]$, $e$ is the charge of the positron, $Q_{\tau}$ is the charge of $\tau$ lepton in units of $e$, $A(s)=F_{2}(q^{2})$ is the Pauli form-factor, and $B(s)=F_{3}(q^{2})$ is the electric dipole form-factor, which depend on $s=q^{2}$, where $q=k_{1}+k_{2}=p_{-}+p_{+}$. At the on-shell photon point, $A(0)$ is an anomalous magnetic dipole moment $a$, while $B(0)$ is related to the $CP$ violating electric dipole moment $d$,
$A(0)=a=\frac{1}{2}(g-2),\qquad\quad B(0)=\frac{2m_{\tau}}{e{Q_{\tau}}}d,$ (4)
where $g$ is the gyromagnetic factor.
@@
$\Gamma^{\mu}_{Z}(q)={{-i}}\frac{g_{Z}}{2}\,\Big{\{}\gamma^{\mu}(v_{\tau}-\gamma_{5}a_{\tau})+\frac{\sigma^{\mu\nu}q_{\nu}}{2m_{\tau}}\,\big{[}iX(s)+Y(s)\gamma_{5}\big{]}\Big{\}},$ (7)
where $g_{Z}=e/(s_{W}c_{W})$ = $2M_{Z}(\sqrt{2}G_{F})^{1/2}$, $s_{W}\equiv\sin\theta_{W}$, $c_{W}\equiv\cos\theta_{W}$; $\theta_{W}$ is the weak mixing angle and $G_{F}=1.1663788(6)\times 10^{-5}$ GeV-2 [20] is the Fermi constant. The vector and axial-vector couplings for the $\tau$ lepton are $v_{\tau}=-1/2+2s_{W}^{2}$ and $a_{\tau}=-1/2$, respectively, $X(s)$ is the weak anomalous magnetic form-factor, and $Y(s)$ is related to the $CP$ violating weak electric form-factor [^2] Sometimes these form-factors are defined with the additional factor $(2c_{W}s_{W})^{-1}$ [30, 22, 19]..
@@
$u$ $+\frac{2}{3}$ $+\frac{1}{2}-\frac{4}{3}s_{W}^{2}$ $+\frac{1}{2}$
@@
$d$ $-\frac{1}{3}$ $-\frac{1}{2}+\frac{2}{3}s_{W}^{2}$ $-\frac{1}{2}$
@@
To simplify calculations, it is convenient to use the Gordon identities for the matrix elements of $\Gamma_{\gamma}^{\mu}(q)$ and $\Gamma_{Z}^{\mu}(q)$. This allows one to rewrite the respective currents as
$\bar{u}(p_{-})\,\Gamma^{\mu}_{\gamma}(q)\,v(p_{+})$ $=$ $-ie{{Q_{\tau}}}\,\bar{u}(p_{-})\Big{\{}\gamma^{\mu}(1+A(s))$ (8)
$+\frac{(p_{+}-p_{-})^{\mu}}{2m_{\tau}}\big{[}A(s)-iB(s)\gamma_{5}\big{]}\Big{\}}v(p_{+}),$
$\bar{u}(p_{-})\,\Gamma^{\mu}_{Z}(q)\,v(p_{+})$ $=$ ${{-i}}\frac{g_{Z}}{2}\,\bar{u}(p_{-})\Big{\{}\gamma^{\mu}((v_{\tau}+X(s))-\gamma_{5}a_{\tau})$ (9)
$+\frac{(p_{+}-p_{-})^{\mu}}{2m_{\tau}}\big{[}X(s)-iY(s)\gamma_{5}\big{]}\Big{\}}v(p_{+}).$
Next we introduce radiative corrections following Refs. [25, 17]. The amplitude, called ${\cal M}^{IBA}$, for arbitrary initial ($i$) and final ($f$) fermions can be written as
${\cal M}^{IBA}$ $=$ $\frac{e^{2}Q_{f}Q_{i}}{s}V_{fi}(s,t)\,\gamma_{\mu}\otimes\gamma^{\mu}$
$+\Bigl{(}\frac{g_{Z}}{2}\Bigr{)}^{2}\frac{Z_{fi}(s,t)}{d(s)}\,\gamma_{\mu}[v_{i}(s,t)-a_{i}\gamma_{5}]\otimes\gamma^{\mu}[v_{f}(s,t)-a_{i}\gamma_{5}],$
where the following shortcut notation is used: operator on the left of $\otimes$ is sandwiched between the spinors $\bar{v}(k_{2})\ldots u(k_{1})$, and operator on the right of $\otimes$ is sandwiched between the spinors $\bar{u}(p_{-})\ldots v(p_{+})$. In Eq. (2.1) $Q_{i}$ and $Q_{f}$ are the charges of initial and final fermions in units of $e$.
Furthermore, $t=(k_{1}-p_{-})^{2}$, $d(s)=s-M_{Z}^{2}+is\,\Gamma_{Z}/M_{Z}$ with running $Z$-boson decay width and $Z_{fi}(s,t)$ is the normalization correction for the $Z$-boson propagator. $v_{i}(s,t)$ and $v_{f}(s,t)$ are modified vector couplings
$v_{i}(s,t)=T_{3i}-2Q_{i}s_{W}^{2}K_{i}(s,t),\qquad v_{f}(s,t)=T_{3f}-2Q_{f}s_{W}^{2}K_{f}(s,t),$ (11)
where $T_{3}$ is the 3rd component of the fermion weak isospin. In Eq. (2.1), the factor $V_{fi}(s,t)$ includes the terms:
$V_{fi}(s,t)=\Gamma_{vp}(s)+\Bigl{(}\frac{g_{Z}}{e}\Bigr{)}^{2}s_{W}^{4}Z_{fi}(s,t)\frac{s}{d(s)}[K_{fi}(s,t)-K_{f}(s,t)K_{i}(s,t)],$ (12)
where $\Gamma_{vp}(s)$ includes re-summed vacuum-polarization loop contributions to the photon propagator, and the complex EW form-factors $K_{i}(s,t)$, $K_{f}(s,t)$ and $K_{fi}(s,t)$ are defined in Ref. [17].
Eq. (2.1) represents improved photon exchange amplitude with running QED constant and improved $Z$-boson exchange which include:
  • corrections to the photon propagator coming from the vacuum-polarization loops,
  • corrections to the $Z$-boson propagator and couplings embedded in the form-factors $Z_{fi}(s,t)$, $K_{i}(s,t),\,K_{f}(s,t)$ and $K_{fi}(s,t)$,
  • contributions from the $WW$ and $ZZ$ box diagrams also included in the form-factors,
  • mixed ${\cal O}(\alpha\alpha_{s},\,\alpha\alpha_{s}^{2},\ldots)$ corrections originating from gluon insertions in the self-energy loop diagrams.
The EW form-factor corrections are available in Dizet library [25]. For further details we refer to Ref. [17]. Note that corrections in IBA become numerically sizable at high-energies, well above $Z$-boson peak. At the lower energies one can use the amplitude in which all the form-factors and $\Gamma_{vp}(s)$ are replaced by 1.
Next we add the amplitude due to the dipole form-factors of the $\tau$ lepton, introduced in (3) and (7). Using identities (8) and (9), one obtains
${\cal M}^{DM}=\frac{e^{2}Q_{f}Q_{i}}{s}V_{fi}(s,t)\,\gamma_{\mu}\otimes[A\gamma^{\mu}+\frac{(p_{+}-p_{-})^{\mu}}{2m_{\tau}}(A-iB\gamma_{5})]$ (13)
$+\Bigl{(}\frac{g_{Z}}{2}\Bigr{)}^{2}\frac{Z_{fi}(s,t)}{d(s)}\,\gamma_{\mu}[v_{i}(s,t)-a_{i}\gamma_{5}]\otimes[X\gamma^{\mu}+\frac{(p_{+}-p_{-})^{\mu}}{2m_{\tau}}(X-iY\gamma_{5})],$
and the total amplitude is ${\cal M}={\cal M}^{IBA}+{\cal M}^{DM}$.
Spin-correlation matrix
@@
If $\tau$ decays are taken into account, the vectors defining $\tau^{-}$ and $\tau^{+}$ density matrices, $s^{-}_{i}$ and $s^{+}_{j}$ in Eq. (18), are replaced respectively by the polarimetric vectors, depending on the $\tau$-decay matrix elements, $h^{-}_{i}$ and $h^{+}_{j}$.
@@
where $R^{(\gamma)}_{i,j}$, $R^{(Z)}_{i,j}$ and $R^{(\gamma Z)}_{i,j}$ represent contributions due to $\gamma$ exchange, $Z$-boson exchange and $\gamma\,Z$ interference, respectively. In the following, we include only terms linear in the dipole form-factors. The quantization of $\tau^{\mp}$ are performed in their respective rest frames. These frames differ from the reaction frame, defined by Eq. (2), by the boosts along $\tau^{\mp}$ directions.
For energies below and around $Z$-boson peak, we show results with radiative corrections switched off, as they can be there safely neglected (incorporated into redefinition of constants). For higher energies, the phenomenological picture would be far more complicated also because of experimental detection/reconstruction criteria, and possibly initial state bremsstrahlung photons of $p_{T}\sim m_{\tau}$ getting lost in the beam-pipe. That is why, this regime remain out of scope of the present work [^3] Note that in contrary to Belle II energies or c.m. energy corresponding to the Z peak, at higher energies presence of initial state photons of $p_{T}$ comparable or higher than the tau mass and at the same time lost in the beam-pipe are not disfavored. This complicates construction of useful for our purposes observables. At the Z-peak energies, such photons are disfavored because then amplitudes are not resonant. At low energies, the ratio of tau mass to beam energy is much larger than the detector minimal acceptance angle. .
@@
$R_{11}^{(Z)}=\frac{g_{Z}^{4}m_{\tau}^{4}\gamma^{2}}{4N_{i}D_{Z}(s)}\sin^{2}(\theta)\big{\{}\left(1-\gamma^{2}\right)a_{\tau}^{2}+v_{\tau}\bigl{[}4\gamma^{2}\text{Re}(X(s))$
@@
$R_{12}^{(Z)}=-\frac{g_{Z}^{4}m_{\tau}^{4}\beta\gamma^{4}}{2N_{i}D_{Z}(s)}{\sin^{2}(\theta)\big{[}\text{Im}(X(s))a_{\tau}-\text{Re}(Y(s))v_{\tau}\big{]}\left(a_{i}^{2}+v_{i}^{2}\right)},$
$R_{22}^{(Z)}=\frac{g_{Z}^{4}m_{\tau}^{4}\gamma^{2}}{4N_{i}D_{Z}(s)}\left(\gamma^{2}-1\right){\sin}^{2}(\theta)\left(a_{\tau}^{2}-v_{\tau}^{2}\right)\left(a_{i}^{2}+v_{i}^{2}\right),$
$R_{13}^{(Z)}=\frac{g_{Z}^{4}m_{\tau}^{4}\gamma^{3}}{4N_{i}D_{Z}(s)}\big{\{}4\,\beta\,\text{sin}(\theta)\,a_{i}\left[\gamma^{2}\text{Re}(X(s))a_{\tau}+\left(-\gamma^{2}\text{Im}(Y(s))+a_{\tau}\right)v_{\tau}\right]v_{i}$
@@
$R_{23}^{(Z)}=-\frac{g_{Z}^{4}m_{\tau}^{4}\gamma^{3}}{2N_{i}D_{Z}(s)}\text{sin}(\theta)\big{\{}\beta\gamma^{2}\text{Re}(Y(s))\left[v_{\tau}\text{cos}(\theta)(a_{i}^{2}+v_{i}^{2})+2\beta a_{\tau}a_{i}v_{i}\right]$